Session: Feb 2011 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (orfei at Sat Feb 26 19:04:24 2011) Minor failures:
MK5 failure at 17:48. Restart schedule at 18:14, line 500.

Onsala (Michael Lindqvist at Sun Feb 27 09:43:58 2011) Minor failures:
Communication problems between FS and VLBI rack (see ERRORs in the log). Not sure why.

Westerbork (Geert Kuper at Mon Feb 28 07:33:33 2011) Minor failures: SCHED
First scan for wb (scan 8) byond HA limits

Torun (Kaz Borkowski at Mon Feb 28 08:40:10 2011) Minor failures: POIN FS
FS lost connection with telescope control system at 5:01. Observations resumed after about 0.5 hour. Then there were troubles with the telescope control between 9 and 10 UT (some 10 minutes could be lost). Also slower telescope movements (due to low temperature) could affect some pointings to new sources.

Effelsberg (Uwe Bach at Tue Mar 01 14:31:41 2011) Minor failures:
Big problems to find proper attenuation. Cyg A is far too bright

Jodrell_Bank (Alastair Gunn at Tue Mar 29 11:26:43 2011) Minor failures: ANT
1 minute lost at 05:00 and 07:33 due to control problem. VCs 1 and 2 out of Rx band - set to duplicate 3 and 4. No other known problems.